A person incurs a 10 % loss by selling a watch for Rs. 1800. At what price should the watch be sold to earn a 10 % profit?
 
Friends…!
Let the cost price of the watch be ₹ 100
The person incurs a loss of 10% by selling a watch, it means
The selling price of watch on cost price 100 is
= 100 - 10
=₹ 90.
For 10% profit, the selling price has to be
100 + 10 = ₹ 110.
Thus, the selling price has to be changed from ₹90 to ₹110 on watch cost of Rs. 100.
On a selling price of Rs. 2280, the new selling price with 10% profit has to be
= 2280 × 110 / 90
= 180 × 110/9
= 20×11
= 2200
 

Grade:12

2 Answers

Arun
4 years ago

Let the new S.P be x, then

(100 - loss%) : (1st S.P.) = (100 + gain%) : (2nd S.P.)

 

90/1800 = 110/x

 

x = 110*20 = 2200 Rs

Vikas TU
4 years ago
 
Dear student 
Please do not do stupid things here , You are here for ask questions. 
Hope you will understand 
Good Luck